OpenStudy (anonymous):

Solve the exponential equation 3^x^2+5x=1/81

OpenStudy (anonymous):

\[3^{x ^{2}+5x}=\frac{ 1 }{81 }=\frac{ 1 }{3^{4} }=3^{-4}\] \[x ^{2}+5x+4=0\] make factors and find the values of x

OpenStudy (anonymous):

(x-5)(x+1) x=5 x=-1

OpenStudy (anonymous):

wrong.

OpenStudy (anonymous):

\[(x+1)(x+4)=0~~~~~~~~~~~~~~~~~~~~~~~~x=-1,~~~~~~x=-4\]